Elif can produce 6 pies or 30 cakes in 1 hour. Ahmet can produce 10 pies or 20 cakes in 1 hour.c) What is the opportunity cost of producing pies for Elif? What is the opportunity cost of producing pies for Ahmet? d) What is the opportunity cost of producing cakes for Elif? What is the opportunity cost of producing cakes for Ahmet? e) Who has a comparative advantage in producing pies and who has a comparative advantage in producing cakes? f) On the PPF lines, show what Elif produces and what Ahmet produces when they specialize. g) Suppose that they set the exchange price as 1 Pie = 3 Cakes. If they specialize and trade, show an exchange situation that is beneficial for both (compared to the situation you found at part (b))? h) What is the minimum and maximum exchange price they could have determined for an exchange that will be beneficial for both of them?



1
Expert's answer
2021-12-19T18:09:31-0500

c) The opportunity cost of producing pies for Elif is: 30/6 = 5 cakes.

The opportunity cost of producing pies for Ahmet is: 20/10 = 2 cakes.

d) The opportunity cost of producing cakes for Elif is: 6/30 = 0.2 pies.

The opportunity cost of producing cakes for Ahmet is: 10/20 = 0.5 pies.

e) Ahmet has a comparative advantage in producing pies and Elif has a comparative advantage in producing cakes.

f) Elif produces only cakes, point (30; 0) on the PPF, and Ahmet produces only pies, point (0; 10) on the PPF when they specialize.

g) If they set the exchange price as 1 Pie = 3 Cakes, and if they specialize and trade, then an exchange situation is beneficial for both, because both can produce more of both goods.

h) The minimum and maximum exchange prices they could have determined for an exchange that will be beneficial for both of them are:

Between 2 and 5 cakes for one pie and between 0.2 and 0.5 pies for one cake.